# Joint Processing and Transmission Energy Optimization for ISAC in Cell-Free Massive MIMO with URLLC

### v1

@@ VI Joint blocklength and Power Optimization
In this paper, we aim to optimize the blocklength and the power control coefficients to minimize the energy consumption while URLLC and sensing requirements are satisfied. In the following subsections, we cast two non-convex optimization problems: i) minimizing only the energy consumption for transmission; ii) minimizing E2E energy consumption by taking into account the energy consumption at the radio site and the cloud. [^2] In addition to considering radio and cloud processing energy consumption, it is important to take fronthaul energy consumption into account when assessing E2E energy consumption. However, in our analysis, where the primary focus is on the interplay between processing and radio resources, we treat fronthaul energy consumption as a fixed component and do not include it in our considerations.
@@ VI-A Minimizing Energy Consumption for Transmission
$P_{k}\leq P_{\rm tx},\quad\quad\textrm{for}\quad k=1,\cdots,N_{\rm tx}$ (63e)
$P_{\rm tr}=\sum_{k=1}^{N_{\rm tx}}P_{k}$ is the average power consumption for transmission, given as $P_{\rm tr}=\sum_{j=0}^{N_{\rm ue}}\rho_{j}=\|\boldsymbol{\rho}\|^{2}$, where $\boldsymbol{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$, due to the unit-power centralized precoding vectors. Constraints (63b) and (63c) correspond to the URLLC requirements. The $\gamma_{\rm s}$ is the required sensing SINR that is selected according to the target detection performance requirement and $P_{\rm tx}$ is the maximum transmit power per AP.
@@
Consider the problem in (VI-A) and define the collective vectors $\textbf{r}=[r_{1}\ \ldots\ r_{N_{\rm ue}}]^{T}$ and $\boldsymbol{\chi}=[\chi_{1}\ \ldots\ \chi_{N_{\rm ue}}]^{T}$. The optimal solution $\{\boldsymbol{\rho}^{\star},L^{\star}\}$ of the problem given below is also an optimal solution to (VI-A):
$\underset{\boldsymbol{\rho},\boldsymbol{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0}{\textrm{minimize}}\quad\frac{\|\boldsymbol{\rho}\|^{2}}{\overline{L}}\,$ (64a)
@@
$\left\|\begin{bmatrix}\sqrt{2}\mathsf{a}_{i,0}\sqrt{\rho_{0}}\\ \vdots\\ \sqrt{2}\mathsf{a}_{i,N_{\rm ue}}\sqrt{\rho_{N_{\rm ue}}}\\ \sqrt{2}\mathsf{b}_{i}\sqrt{\rho_{i}}\\ \sqrt{2}\sigma_{n}\\ 1+\chi_{i}\\ r_{i}\end{bmatrix}\!\right\|\leq 1+\chi_{i}+r_{i},\forall i,$ (64k)
$\boldsymbol{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\boldsymbol{\rho}\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}$ (64l)
$\left\|\textbf{F}_{k}\boldsymbol{\rho}\right\|\leq\sqrt{P_{\rm tx}},\quad k=1,\ldots,N_{\rm tx},$ (64m)
$L\leq L_{\rm max},$ (64n)
@@
The optimization problem in (1) is still not convex due to the non-convex constraints (64b), (64c) and (64l). The terms that destroy convexity are the convex terms $1/\overline{L}$ and $\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}$ (in terms of $\boldsymbol{\rho}$ and r) on the right-hand side of (64b) and the left-hand side of (64c), respectively. To this end, we apply the concave-convex procedure (CCP) approach to (64b) and (64c), and the FPP-SCA method [34] to (64l), wherein $-\boldsymbol{\rho}^{T}M\,\textbf{A}_{D}\boldsymbol{\rho}$ is a concave function. Moreover, to avoid any potential infeasibility issue regarding (64l) during the initial iterations of the algorithm, we add slack variable $\chi_{0}\geq 0$ and a slack penalty $\lambda$, to the convexified problem at the initial iterations. In subsequent iterations, we set $\chi_{0}$ to zero if it is less than a threshold, denoted as $\chi_{0}\leq\epsilon_{\chi}$. Finally, the convex problem that is solved at the $c^{\rm th}$ iteration becomes
$\underset{\boldsymbol{\rho},\boldsymbol{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\frac{\|\boldsymbol{\rho}\|^{2}}{\overline{L}}+\lambda\chi_{0}\,,$ (65a)
@@
$\gamma_{\rm s}\boldsymbol{\rho}^{T}\textbf{B}_{D}\boldsymbol{\rho}-2M\,\Re\left(\left(\boldsymbol{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\boldsymbol{\rho}\right)$
$\hskip 14.22636pt\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}-M\left(\boldsymbol{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\boldsymbol{\rho}^{(c-1)}+\chi_{0},$ (65d)
@@
1:Initialization: Set arbitrary $\boldsymbol{\rho}^{(0)},\textbf{r}^{(0)}\geq\textbf{0}$, the solution accuracy parameters $\epsilon,\epsilon_{\chi}>0$, $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, and $\lambda>0$. Set the iteration counter to $c=0$, and a maximum iteration number $c_{\rm max}$. Moreover, let $\mathsf{F}^{(0)}=\infty$ be the initial value of objective function and $\Delta\mathsf{F}^{(c)}=\mathsf{F}^{(c)}-\mathsf{F}^{(c-1)}$ with $\mathsf{F}^{(-1)}=0$.
@@
4:Set $\boldsymbol{\rho}^{(c)}$, $\overline{L}^{(c)}$, and $\textbf{r}^{(c)}$ to the solution of the convex problem in (VI-A), where the previous iterates $\boldsymbol{\rho}^{(c-1)}$, $\overline{L}^{(c-1)}$, and $\textbf{r}^{(c-1)}$ are taken as constant.
@@
7:Output: The transmit power coefficients $\boldsymbol{\rho}^{(c)}$ and the blocklength $L^{(c)}$.
@@ VI-B Minimizing E2E Energy Consumption
$L\leq L_{\rm max},\quad$ (66c)
$\overline{\mathsf{SINR}}_{s}\geq\gamma_{\rm s},\quad$ (66d)
$P_{k}\leq P_{\rm tx},\quad\quad\textrm{for}\quad k=1,\cdots,N_{\rm tx}.$ (66e)
@@
$P_{\rm total}=P_{\rm FIXED}+\Delta^{\rm tr}\|\boldsymbol{\rho}\|^{2}+\frac{1}{L}f_{1}+\frac{L_{d}}{L}f_{2}$ (67)
@@
$\mathsf{F_{2}}\triangleq\frac{1}{B}\left(L\left(P_{\rm FIXED}\right)+\left(L-L_{p}\right)f_{2}+\Delta^{\rm tr}\frac{\|\boldsymbol{\rho}\|^{2}}{\bar{L}}\right).$ (71)
The non-convex optimization problem in (VI-B) can be solved using the same approach in P1 as follows:
$\underset{\boldsymbol{\rho},\boldsymbol{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\mathsf{F_{2}}+\lambda\chi_{0},$ (72a)
@@ VII Numerical Results
In this section, we present numerical results to evaluate the performance of the proposed joint blocklength and power allocation algorithms. We consider a total area of 500 $\rm m\times$ 500 $\rm m$. The sensing location is in the center of the area, i.e., $(250,250)$. The number of antenna elements per AP is set to $M=4$. The $N_{\rm tx}=16$ ISAC transmit APs are uniformly distributed in the area. We consider the number of sensing receive APs equal to $N_{\rm rx}=1$ or $N_{\rm rx}=2$ where the first AP is located at $(200,250)$ and the second one is located at $(300,250)$. The total number of URLLC UEs in the network is $N_{\rm ue}=8$, unless otherwise stated. We set $P_{\rm tx}=100$ mW and the uplink pilot transmission power for each UE is $50$ mW.
The large-scale fading coefficients, shadowing parameters, probability of LOS, and the Rician factors are simulated based on the 3GPP Urban Microcell model, defined in [35, Table B.1.2.1-1, Table B.1.2.1-2, Table B.1.2.2.1-4]. The path losses for the Rayleigh fading target-free channels are also modeled by the 3GPP Urban Microcell model with the difference that the channel gains are multiplied by an additional scaling parameter equal to $0.3$ to suppress the known parts of the target-free channels due to LOS and permanent obstacles [25]. The sensing channel gains are computed by the bi-static radar range equation [26]. The carrier frequency, the bandwidth, and the noise variance are set to $1.9$ GHz, $200$ KHz, and $-114$ dBm, respectively. The number of pilot symbols is $L_{p}=10$.
@@
The spatial correlation matrices for the communication channels are generated by using the local scattering model in [27, Sec. 2.5.3]. The RCS of the target is modeled by the Swerling-I model with $\sigma_{\rm rcs}=0\,\rm dBsm$. For all the UEs, we set the same packet size, the maximum transmission delay, and the DEP threshold to $b_{i}=256$ bits, $D_{i}^{\rm(th)}=1$ ms, and $\epsilon_{\rm th}=10^{-5}$, respectively unless otherwise stated. In the proposed algorithms, $\epsilon=10^{-3}$, $\epsilon_{\chi}=10^{-6}$, $\lambda=10$, and $c_{\rm max}=30$. The sensing SINR threshold is set to $\gamma_{s}=0$ dB, unless otherwise stated. The remaining parameters are detailed in Table II, where the values are consistent with those in [14], except for a scaling of the total processing capacity and the corresponding power of a DU by a factor of ten to accommodate all the processing in a single DU.
We compare the performance of two proposed algorithms: i) SeURLLC+, which aims to minimize the transmission energy consumption for as ISAC cell-free massive MIMO system with URLLC UEs, and ii) E2E SeURLLC+, which aims to minimize the E2E energy consumption. We also consider two baseline algorithms for only URLLC systems without integrating sensing: i) URLLC-only, and ii) E2E URLLC-only, which represent minimizing transmission energy and E2E energy, respectively.
@@
Figure 2: E2E energy consumption over one transmission with different algorithms and with $N_{\rm rx}=1,2$.
In Fig. 2, we illustrate E2E energy consumption with varying algorithms and two values for the number of sensing receive APs. The chart breaks down energy consumption by system components and different operations, including transmission, ISAC APs, sensing receive APs, communication and sensing processing, and “Others” representing load-independent and idle mode power consumption at the cloud. It is shown that in URLLC-only scenarios, the majority of the energy consumption is related to the ISAC APs and then the load-independent and idle mode power consumption at the cloud. Remarkably, E2E energy minimization demonstrates a substantial reduction in energy consumption by approximately $50\%$ in URLLC-only scenarios. This underscores the importance of E2E energy-aware optimization. Conversely, in ISAC scenarios, the impact of E2E energy minimization is less pronounced, resulting in only a marginal decrease. However, we observe that the energy consumption due to sensing processing becomes predominant when $N_{\rm rx}=2$. This observation suggests a potential optimization avenue, as turning off one sensing receive AP can halve the energy consumption, contributing significantly to overall energy efficiency.
(a) E2E URLLC-only
(b) E2E SeURLLC+ , $N_{\rm rx}=1$
(c) E2E SeURLLC+ , $N_{\rm rx}=2$
Figure 3: Energy consumption in one transmission with a) E2E URLLC-only, b) E2E SeURLLC+ with $N_{\rm rx}=1$, and E2E SeURLLC+ with $N_{\rm rx}=2$.
To complete the findings in Fig. 2, we provide a comparative analysis of the energy consumption distribution across various components within the system in Fig. 3. E2E energy minimization algorithms are considered for a) URLLC-only scenario, b) ISAC scenario with 1 sensing receive AP, and c) ISAC scenario with 2 sensing receive APs. As shown in Fig. 2(a) and Fig. 2(b), the ISAC APs are the most energy consuming components. Notably, the increment in the number of sensing receive APs is accompanied by a significant rise in energy consumption corresponding to the processing of the received signals for target detection. This is evident that with the addition of another receive AP, the share of sensing processing increases up to $52\%$, approximately four times larger than the case with $N_{\rm rx}=1$.
Figure 4: Blocklength vs. decoding error probability threshold.
Figure 5: E2E energy consumption vs. decoding error probability threshold.
VII-A Impact of Reliability and Delay Requirements
In this subsection, we investigate the impact of the reliability and delay requirements on the performance of the above-mentioned algorithms. Fig. 4 shows the blocklength as a function of decoding error probability (DEP) threshold. Generally, the blocklength decreases by relaxing the reliability threshold in all the algorithms except URLLC-only. It is shown that the URLLC-only algorithm always ends up with the maximum available blocklength, i.e., $L_{\rm max}$, since increasing the blocklength allows lower transmit power to satisfy the reliability constraint. We also observe that minimizing only the transmission energy consumption will lead to higher blocklength values compared to the E2E energy minimization algorithms, i.e., E2E URLLC-only and E2E SeURLLC+. Compared to URLLC-only, the E2E URLLC-only algorithm results in a significant decrease in the blocklength to minimize the E2E energy consumption over one transmission. For ISAC scenarios, we see the same behaviours as in E2E URLLC-only. However, the blocklength is slightly higher due to the interplay between sensing power consumption and the blocklength as transmit power consumption will decrease with larger blocklength.
Fig. 5 illustrates the E2E energy consumption in terms of the DEP threshold. The results show that relaxing the reliability constraint would generally reduce the E2E energy consumption. However, ISAC scenarios demand higher energy consumption as it is expected. Moreover, we observe that algorithms that minimize E2E energy consumption outperform the corresponding algorithms which only minimize the transmission energy consumption. Particularly, when we do not integrate sensing in the system, there is a noticeable gap between performance of URLLC-only and E2E URLLC-only algorithms.
Figure 6: E2E energy consumption vs. maximum delay threshold.
The results for E2E energy consumption versus maximum delay threshold, $D_{i}^{(\rm tx)}$, are depicted in Fig. 6. As it is shown, E2E energy consumption does not change significantly with different maximum delay threshold, except for URLLC-only algorithm. The reason is that the URLLC-only algorithm would always result in the highest possible blocklength to minimize the transmission energy consumption while the optimized value for the blocklength with other algorithms would be much lower than the one with URLLC-only to minimize the energy consumption. That is why the SeURLLC+ has lower E2E energy consumption than URLLC-only. However, the results of E2E URLLC-only algorithm show that minimum E2E energy consumption could be archived by optimizing the blocklength with respect to E2E energy consumption. This behavior highlights the importance of the E2E energy minimization.

### v2

@@ VI Joint blocklength and Power Optimization
In this paper, we aim to optimize the blocklength and the power control coefficients to minimize the energy consumption while URLLC and sensing requirements are satisfied. In the following subsections, we cast two non-convex optimization problems: i) minimizing only the energy consumption for transmission; ii) minimizing E2E energy consumption by taking into account the energy consumption at the radio site and the cloud. [^2] In addition to considering radio and cloud processing energy consumption, it is important to take fronthaul energy consumption into account when assessing E2E energy consumption. However, in our analysis, where the primary focus is on the interplay between processing and radio resources, we treat fronthaul energy consumption as a fixed component and do not include it in our considerations. The corresponding algorithms to these optimization problems are represented by SeURLLC+ and E2E SeURLLC+, respectively.
@@ VI-A Minimizing Energy Consumption for Transmission
$P_{k}\leq P_{\rm tx},\quad\quad k=1,\cdots,N_{\rm tx}$ (51e)
$P_{\rm tr}=\sum_{k=1}^{N_{\rm tx}}P_{k}$ is the average power consumption for transmission, given as $P_{\rm tr}=\sum_{j=0}^{N_{\rm ue}}\rho_{j}=\|\bm{\rho}\|^{2}$, where $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$, due to the unit-power centralized precoding vectors. Constraints (51b) and (51c) correspond to the URLLC requirements. The $\gamma_{\rm s}$ is the required sensing SINR that is selected according to the target detection performance requirement and $P_{\rm tx}$ is the maximum transmit power per AP.
@@
Consider the problem in (51) and define the collective vectors $\textbf{r}=[r_{1}\ \ldots\ r_{N_{\rm ue}}]^{T}$ and $\bm{\chi}=[\chi_{1}\ \ldots\ \chi_{N_{\rm ue}}]^{T}$. The optimal solution $\{\bm{\rho}^{\star},L^{\star}\}$ of the problem given below is also an optimal solution to (51):
$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0}{\textrm{minimize}}\quad\frac{\|\bm{\rho}\|^{2}}{\overline{L}}\,$ (52a)
@@
$\left\|\begin{bmatrix}\sqrt{2}\mathsf{a}_{i,0}\sqrt{\rho_{0}}\\ \vdots\\ \sqrt{2}\mathsf{a}_{i,N_{\rm ue}}\sqrt{\rho_{N_{\rm ue}}}\\ \sqrt{2}\mathsf{b}_{i}\sqrt{\rho_{i}}\\ \sqrt{2}\sigma_{n}\\ 1+\chi_{i}\\ r_{i}\end{bmatrix}\!\right\|\leq 1+\chi_{i}+r_{i},\forall i,$ (52d)
$\bm{\rho}^{T}\left(\gamma_{\rm s}\textbf{B}_{D}-M\,\textbf{A}_{D}\,\right)\bm{\rho}\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}$ (52e)
$\left\|\textbf{F}_{k}\bm{\rho}\right\|\leq\sqrt{P_{\rm tx}},\quad k=1,\ldots,N_{\rm tx},$ (52f)
$L\leq L_{\rm max},$ (52g)
@@
The optimization problem in (52) is still not convex due to the non-convex constraints (52b), (52c) and (52e). The terms that destroy convexity are the convex terms $1/\overline{L}$ and $\frac{\rho_{i}\mathsf{b}_{i}^{2}}{r_{i}}$ (in terms of $\bm{\rho}$ and r) on the right-hand side of (52b) and the left-hand side of (52c), respectively. To this end, we apply the concave-convex procedure (CCP) approach to (52b) and (52c), and the FPP-SCA method [42] to (52e), wherein $-\bm{\rho}^{T}M\,\textbf{A}_{D}\bm{\rho}$ is a concave function. Moreover, to avoid any potential infeasibility issue regarding (52e) during the initial iterations of the algorithm, we add slack variable $\chi_{0}\geq 0$ and a slack penalty $\lambda$, to the convexified problem at the initial iterations. In subsequent iterations, we set $\chi_{0}$ to zero if it is less than a threshold, denoted as $\chi_{0}\leq\epsilon_{\chi}$. Finally, the convex problem that is solved at the $c^{\rm th}$ iteration becomes
$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\frac{\|\bm{\rho}\|^{2}}{\overline{L}}+\lambda\chi_{0}\,,$ (53a)
@@
$\gamma_{\rm s}\bm{\rho}^{T}\textbf{B}_{D}\bm{\rho}-2M\,\Re\left(\left(\bm{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\bm{\rho}\right)$
$\hskip 14.22636pt\leq-\gamma_{\rm s}MN_{\rm rx}\sigma_{n}^{2}-M\left(\bm{\rho}^{(c-1)}\right)^{T}\textbf{A}_{D}\bm{\rho}^{(c-1)}+\chi_{0},$ (53d)
@@
1:Initialization: Set arbitrary $\bm{\rho}^{(0)},\textbf{r}^{(0)}\geq\textbf{0}$, the solution accuracy parameters $\epsilon,\epsilon_{\chi}>0$, $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, and $\lambda>0$. Set the iteration counter to $c=0$, and a maximum iteration number $c_{\rm max}$. Moreover, let $\mathsf{F}^{(0)}=\infty$ be the initial value of objective function and $\Delta\mathsf{F}^{(c)}=\mathsf{F}^{(c)}-\mathsf{F}^{(c-1)}$ with $\mathsf{F}^{(-1)}=0$.
@@
4:Set $\bm{\rho}^{(c)}$, $\overline{L}^{(c)}$, and $\textbf{r}^{(c)}$ to the solution of the convex problem in (53), where the previous iterates $\bm{\rho}^{(c-1)}$, $\overline{L}^{(c-1)}$, and $\textbf{r}^{(c-1)}$ are taken as constant.
@@
7:Output: The transmit power coefficients $\bm{\rho}^{(c)}$ and the blocklength $L^{(c)}$.
@@ VI-B Minimizing E2E Energy Consumption
[…]
@@
$P_{\rm total}=P_{\rm FIXED}+\Delta^{\rm tr}\|\bm{\rho}\|^{2}+\frac{1}{L}f_{1}+\frac{L_{d}}{L}f_{2}$ (55)
@@
$\mathsf{F_{2}}\triangleq\frac{1}{B}\left(L\left(P_{\rm FIXED}\right)+\left(L-L_{p}\right)f_{2}+\Delta^{\rm tr}\frac{\|\bm{\rho}\|^{2}}{\bar{L}}\right).$ (59)
The non-convex optimization problem in (54) can be solved using the same approach in P1 as follows:
$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\mathsf{F_{2}}+\lambda\chi_{0},$ (60a)
@@ VII Numerical Results
In this section, we present numerical results to evaluate the performance of the proposed joint blocklength and power allocation algorithms. We consider a total area of 500 $\rm m\times$ 500 $\rm m$, with the sensing location positioned at the center. The $N_{\rm tx}=16$ ISAC transmit APs are uniformly distributed in the area. We consider the number of sensing receive APs equal to $N_{\rm rx}=1$ or $N_{\rm rx}=2$ where the first AP is located at $(200,250)$ and the second one is located at $(300,250)$. The number of antenna elements per AP is set to $M=4$. The total number of URLLC UEs in the network is $N_{\rm ue}=8$. We set $P_{\rm tx}=100$ mW and the uplink pilot transmission power for each UE is $50$ mW.
The large-scale fading coefficients, shadowing parameters, probability of LOS, and the Rician factors are simulated based on the 3GPP Urban Microcell model, defined in [43, Table B.1.2.1-1, Table B.1.2.1-2, Table B.1.2.2.1-4]. The path losses for the Rayleigh fading target-free channels are also modeled by the 3GPP Urban Microcell model with the difference that the channel gains are multiplied by an additional scaling parameter equal to $0.3$ to suppress the known parts of the target-free channels due to LOS and permanent obstacles [34]. The sensing channel gains are computed by the bi-static radar range equation [9]. The carrier frequency, the bandwidth, and the noise variance are set to $1.9$ GHz, $200$ KHz, and $-114$ dBm, respectively. The number of pilot symbols is $L_{p}=10$.
The spatial correlation matrices for the communication channels are generated by using the local scattering model in [35, Sec. 2.5.3]. The RCS of the target is modeled by the Swerling-I model with $\sigma_{\rm rcs}=0\,\rm dBsm$. For all the UEs, the packet size, maximum transmission delay, and DEP threshold are $b_{i}=256$ bits, $D_{i}^{\rm(th)}=1$ ms, and $\epsilon_{\rm th}=10^{-5}$, respectively. In the proposed algorithms, $\epsilon=10^{-3}$, $\epsilon_{\chi}=10^{-6}$, $\lambda=10$, and $c_{\rm max}=30$. The sensing SINR threshold is set to $\gamma_{s}=0$ dB, unless otherwise stated. The remaining parameters are detailed in Table II, where the values are consistent with those in [30], except for a scaling of the total processing capacity and the corresponding power of a DU by a factor of ten to accommodate all the processing in a single DU.
@@
Figure 2: (a) Transmission energy consumption, (b) E2E energy consumption, and (c) breakdown of energy consumption over one transmission for different optimization algorithms, for clutter-aware target detector with $N_{\rm rx}=2$ and $\gamma_{s}=0$ dB.
We compare the performance of two proposed algorithms: i) SeURLLC+ aiming to minimize the transmission energy consumption, and ii) E2E SeURLLC+ aiming to minimize the E2E energy consumption. Additionally, we include two baseline algorithms for URLLC systems without sensing integration: (i) URLLC-only, which minimizes transmission energy consumption, and (ii) E2E URLLC-only, which minimizes E2E energy consumption. We further compare results using clutter-aware and clutter-unaware MAPRT detectors, as defined in (17) and (18), respectively.
Fig. 2(a) illustrates the transmission energy consumption for four algorithms with the clutter-aware detector. The results show that integrating sensing into the system necessitates higher transmission energy. Notably, the URLLC-only and SeURLLC+ achieve lower transmission energy compared to their E2E counterparts, as their primary objective is to minimize transmission energy consumption.
The E2E energy consumption is depicted in Fig. 2(b). We observe that E2E energy minimization results in a substantial reduction in total energy consumption by approximately $50\%$ in URLLC-only scenarios, while its impact is less pronounced in ISAC scenarios, resulting in only a marginal decrease. The E2E algorithms try to minimize the total energy consumption by reducing the blocklength. However, higher power coefficients are required to meet the reliability requirements for URLLC UEs. In SeURLLC+, higher power coefficients for sensing allow lower blocklength, narrowing the gap in optimal blocklengths between SeURLLC+ and E2E SeURLLC+.
Fig. 2c provides a detailed breakdown of energy consumption by system components and different operations, including ISAC transmit APs, sensing receive APs, communication and sensing processing, and “Others,” which represent load-independent and idle mode power consumption in the cloud. For both algorithms, most of the energy consumption is attributed to sensing processing tasks, ISAC APs, and load-independent and idle mode power consumption in the cloud. However, the E2E SeURLLC+ algorithm slightly reduces the corresponding energy consumption.
Figure 3: (a) Blocklength, (b) E2E energy consumption, (c) detection probability, and (d) sensing update rate vs. decoding error probability threshold with clutter-aware target detector, $N_{\rm rx}=2$ and $\gamma_{s}=0$ dB.
VII-A Impact of URLLC Requirements
Fig. 3 illustrates the blocklength, E2E energy consumption, detection probability and sensing update rate as functions of DEP threshold. Generally, the blocklength in Fig. 3(a) decreases as the reliability threshold is relaxed in all algorithms except for URLLC-only. The URLLC-only algorithm consistently uses the maximum available blocklength, i.e., $L_{\rm max}$. By integrating sensing, the optimal blocklength is significantly reduced compared to URLLC-only, as meeting sensing constraint requires higher power consumption resulting in lower blocklength. Relaxing the reliability constraint would generally reduce the E2E energy consumption. However, ISAC scenarios demand higher energy consumption as it is expected as shown in Fig. 3(b).
Fig. 3(c) and Fig. 3(d) demonstrate the impact of the DEP threshold on sensing performance. Relaxing the DEP constraint slightly reduces detection probability while increasing the sensing update rate, due to smaller blocklengths at higher DEP thresholds. Notably, the E2E SeURLLC+ algorithm outperforms the SeURLLC+ algorithm in both sensing and communication performance.
It is worth mentioning that higher delay threshold allows the system to operate with higher blocklength. However, it does not effect the results since the algorithm aims to minimize the energy consumption. Therefore, it still chooses a smaller blocklength.